OpenStudy (agl202):

Quick Help... Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3. −15 −8 10 25

jimthompson5910 (jim_thompson5910):

g(x) = x-7 g(-3) = -3 - 7 ... replace every x with -3 g(-3) = _______ (fill in the blank)

OpenStudy (agl202):

-10

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so saying f(g(-3)) is the same as saying f(-10) I replaced the g(-3) with -10 since the two are equal

OpenStudy (elonasushchik):

hmm...

jimthompson5910 (jim_thompson5910):

now evaluate f(-10)

OpenStudy (agl202):

I knew it! Just checking! thx...:D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (agl202):

@ElonaSushchik Poshla Von!!!!